We consider the sequence of real numbers (Un) defined on N by Uo = -1, U1 = 1/2 and for every n E N, U(n+2) = U(n+1) - 1/4 Un. Where N represents the set of natural numbers
(i) Calculate Vo.
(ii) Show that (Vn) is a geometric sequence and precise its common ratio.
(iii) Express Vn in terms of n and calculate its limit.
1
Expert's answer
2021-01-29T05:05:59-0500

U(n+2) = U(n+1) - 1/4 Un

Vn = U(n+1) - (1/2)Un.

(i)

"V(0)=U(0+1)-(1\/2)U(0)"

"=1\/2-(1\/2)(-1)=1"

(ii)

"U(0)=-1,"

"U(1)=1\/2,"

"U(2)=1\/2+1\/4=3\/4"

"U(3)=3\/4-1\/8=5\/8"

"U(4)=5\/8-1\/8=7\/16"

"U(5=7\/16-5\/32=9\/32"

"..."

"U(k)=(2k-1)\/2^k"

Prove by mathematical induction

Let the statement "P(n)" given as

"P(n):U(n)=(2n-1)\/2^n"

Then

"P(0):U(0)=(2(0)-1)\/2^0=-1" is true.

"P(1): U(1)=(2(1)-1)\/2^1=1\/2" is true.

Assume that "P(n)" is true for some natural number "k" , i.e.,

"P(k):U(k)=(2k-1)\/2^k"

"P(k+1):U(k+1)=(2(k+1)-1)\/2^{k+1}"

Now, to prove that "P(k+2)" is true, we have


"P(k+2): U(k+2)=U(k+1)-(1\/4)U(k)="

"=\dfrac{2(k+1)-1}{2^{k+1}}-\dfrac{2k-1}{(4)(2^{k})}=\dfrac{4k+4-2-2k+1}{2^{k+2}}"

Thus "P(k+2)" is true, whenever "P(k)" and "P(k+1)" is true. Hence, by the Principle of Mathematical Induction "P(n)" is true for all natural numbers "n"


"U(n)=(2n-1)\/2^n, n=0,1,2,..."

Then


"V(n)=U(n+1)-(1\/2)U(n)"

"=\dfrac{2(n+1)-1}{2^{n+1}}-\dfrac{2n-1}{2^{n+1}}=\dfrac{1}{2^n}, n=0,1,2,..."

Then "(V(n))" is a geometric sequence and its common ratio is "1\/2"


(iii)


"V(n)=\dfrac{1}{2^n}, n=0,1,2,..."

"\lim\limits_{n\to \infin}V(n)=\lim\limits_{n\to \infin}\dfrac{1}{2^n}=0"